OpenStudy (anonymous):

Graph The Function Y=(x-1)^2+3. How Is The Graph A Translation Of f(x)=x^2?

OpenStudy (gorv):

y-3=(x-1)^2 it is shifted parabola

OpenStudy (gorv):

similar to y=x^2 but it is shifted from (0,0) to (1,3)

OpenStudy (anonymous):

Oh Okay Thank You c:

OpenStudy (gorv):

wlcme :)